OpenStudy (anonymous):

Solve the equations (3x+2)(x-4) 0 (x+1)(x-2) = 0 x^2-6x+58=0 I don't get how to solve these, wouldn't x = 0 in all of them?

OpenStudy (azureilai):

No it wouldn't, since there are a second term in each parenthesis where the x is not being multiplied by. For example, the first one if x was zero, it would look like this (3(0)+2)(0-4) 2*-4=-8 The result would not be zero. So in order to find x, you have to factor. What you do is set each term within the parentheses equal to zero and solve 3x+2 =0 x=-2/3 x-4=0 x=4 Do the same for the others

OpenStudy (anonymous):

so for the first on x = 4? How did you get that? I don't understand this

OpenStudy (azureilai):

For the first one, x has two solutions. It is 4 and -2/3. What I did was take the equation inside each parentheses and set them equal to zero. This way, you can get the value of x when y=0

OpenStudy (joannablackwelder):

Azureilai is using the property that states, if quantities are multiplied together equal zero, ether at least one of those quantities must equal 0. Does that help?

OpenStudy (joannablackwelder):

So, the first one could be set up to have x-4=0 or 3x+2=0. You would need to solve both of these equations for x separately.

OpenStudy (anonymous):

fiirst one is 4 second one is 2 third one i have no idea

OpenStudy (joannablackwelder):

No, the first equation could equal 4, but it could also equal -2/3.

OpenStudy (joannablackwelder):

2 is one solution for the second equation but there is another solution for that one too. Solve x+1=0.

OpenStudy (joannablackwelder):

The third one you will need to factor first or use the quadratic formula.

OpenStudy (anonymous):

For the first one isnt 4 he highest soluton? that's what its asking for

OpenStudy (joannablackwelder):

Oh, I didn't realize you were looking for the highest solution. Yes, 4 would be the highest solution.

OpenStudy (anonymous):

same with number 2 then right?

OpenStudy (joannablackwelder):

Yes.

OpenStudy (anonymous):

I don't know how to factor so i don't really know how to do number 3

OpenStudy (anonymous):

and number 2 is 2 right?

OpenStudy (anonymous):

or 4?

OpenStudy (joannablackwelder):

Yes, the highest solution to the second equation is 2.

OpenStudy (joannablackwelder):

Do you know the quadratic formula?

OpenStudy (joannablackwelder):

Since you don't know how to factor, I am assuming you aren't familiar with imaginary numbers. The last one would have no solution unless you can write it with imaginary numbers.